OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. If 3a - 9 = 6, what is the value of 4a - 1?

OpenStudy (anonymous):

3a - 9 = 6 equate for a: 3a = 6 + 9 3a = 12 divide both sides by 3 a = 4

OpenStudy (anonymous):

substitute a to 4a - 1, we get 4(4) - 1 = 15

OpenStudy (anonymous):

is that the answer?!

OpenStudy (anonymous):

yes. and please solve it for you to practice. :)